OpenStudy (moongazer):

The average of X numbers is 5. If Y is added to each number, the average will be?

OpenStudy (anonymous):

average will be 5+Y

OpenStudy (moongazer):

why??

OpenStudy (anonymous):

the sum of X no. was =5X when u add Y to each no. then the sum is 5X+XY=X(5+Y) so the average is x(5+Y)/Y =5+x